OpenStudy (countrygirl1431):

If the equation of a circle is (x + 5)2 + (y - 7)2 = 36, its radius is 6 16 36

OpenStudy (countrygirl1431):

i think its 6

OpenStudy (xapproachesinfinity):

in general circle's equation is (x-a)^2+y-b)^2=r^2 r be the raduis and (a,b) be the center so try to write 36 as square

OpenStudy (xapproachesinfinity):

yes it is :)

OpenStudy (er.mohd.amir):

open bracket and compare ax2+by2+2gx+2fy+c=0 radius=sqrt(g*g+f*f-c)

OpenStudy (xapproachesinfinity):

no need just rewrite 36 as 6^2 and 6 be the radius that's all